OpenStudy (anonymous):

solve the quadratic equation y^2-7y-18=0 by factoring

OpenStudy (anonymous):

x= -2 x=9

OpenStudy (anonymous):

err Y I mean.

OpenStudy (anonymous):

a= 1, b= -7, c= -18 x = [– b ± √b² – 4ac]/2a now u can solve it right

OpenStudy (anonymous):

y ^{2}-9y+2y-18=0 thus y(y-9)+2(y-9)=0 thus (y+2)(y-9)=0 hence y=9 ; x=-2